OpenStudy (anonymous):

Raman's present age is three times his daughter's and nine-thirteenth of his mother's present age. The sum of the present ages of all three of them is 125 years. What is the difference between the present ages of Raman's daughter and Raman's mother?

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

InsanelyChaotic can u help me out this problem

Directrix (directrix):

Let's give this a try.

Directrix (directrix):

Let R's daughter's age be represented by x.

Directrix (directrix):

>Raman's present age is three times his daughter's So, R = 3x because the daughter's age is x and R's age is 3 times that of his daughter.

Directrix (directrix):

> R's age is nine-thirteenth of his mother's R = 9/13 * M where M is R's mother's current age But R = 3x.

Directrix (directrix):

R = 9/13 * M 3x = 9/13 * M Solve for the Mother's age in terms of x.

Directrix (directrix):

3x = 9/13 * M (13/9) * (3x) = (13/9)* (9/13) * M (39/9) x = 1* M (39/9) x = M

Directrix (directrix):

> The sum of the present ages of all three of them is 125 years. -------------------- R + D + M = 125 3x + x + (39/9) x = 125 Solve for x.

Directrix (directrix):

Then, back substitute to get the ages of the daughter and the mother and subtract. @genita

OpenStudy (anonymous):

Solve the following for R, D and M:\[\left\{R=3D,R=\frac{9}{13}M,R+D+M=125\right\}\]\[\{R=45,D=15,M=65\} \]